# OntoURL: A Benchmark for Evaluating Large Language Models on Symbolic Ontological Understanding, Reasoning and Learning

### v2

@@ Appendix D Task Details
Task Definition, Construction Process and Examples
Task U1: Class Definition Understanding
The task is about mapping the correct definition to a given class, which address the ability of LLM to understand a specific class/concept correctly.
Our task generation pipeline processes each ontology file by 1.loading it, extracting all classes with explicit definitions, and then for each class: 2. Retrieves its human-readable label and definition. 3. Selects three distractors, preferably from the same superclass, or randomly if not feasible. 4. Shuffles the correct definition with the distractors to create a four-option set.
Table 6: The instruction and an example of Task U1 Class Definition Understanding.
Instruction: Given a multiple-choice question with options, select the one correct definition and output only the letter.
Question: In the ontology of Plant, which best describes ‘calyx’?
A: A collective phyllome structure that is composed of one or more petals, comprising the inner whorl of non-reproductive floral organs and surrounds the androecium and the gynoecium.
B: A collective phyllome structure composed of all of the carpels in a flower.
C: A collective phyllome structure that encloses a fruit of the Poaceae (caryopsis) and develops from a dried lemma and palea.
D: A collective phyllome structure that is composed of one or more sepals, and comprises the outer whorl of non-reproductive floral organs.
Answer: D
Task U2: Class Relation Understanding
The task focuses on evaluating the ability of LLMs to recognize relational knowledge in ontologies, such as subclass, disjoint, or equivalent class relations.
The question generation pipeline processes each ontology file by: 1. Loading the ontology and extracting all classes along with their hierarchical and relational information. 2. Extracting relation triples (e.g., subclassOf, disjointWith, unionOf) using explicit axioms, object properties, and RDF list structures. 3. Selecting distractors for each relational triple, prioritizing disjoint classes and non-ancestor classes to ensure semantic plausibility. 4. Generating multiple-choice questions, each asking about the correct relational target and pairing it with three distractors (choose from subclass, subsumptions or siblings).
Table 7: The instruction and an example of Task U2 Class Relation Understanding.
Instruction: Given a multiple-choice question with options, select the one correct class and output only its letter.
Question: In the ontology of Human Phenotype, which of the following classes is the superclass of ’regulation of leukocyte differentiation’?
A. regulation of hemopoiesis
B. Parathyroid dysgenesis
C. Olfactory lobe agenesis
D. positive regulation of G protein-coupled receptor signaling pathway
Answer: A
Task U3: Property Domain Understanding
The task targets LLMs’ understanding of data property semantics in ontologies by testing their ability to identify valid domain or range classes.
Each ontology is processed by: 1. Loading the ontology and collecting all data properties and classes. 2. Extracting (property, domain, range) triples for properties with defined domains. 3. Selecting distractors from unrelated classes to ensure plausible options. The distractors are selected by the distance of from the correct class.
Table 8: The instruction and an example of Task U3 Property Domain Understanding.
Instruction: Given a multiple-choice question with options, select the one correct property and output only the letter.
Question: In the ontology of Video Game, which of the following is a valid domain for the data property ’event time’?
A. menu event
B. ProductOrService
C. Character
D. instantaneous event
Answer: D
Task U4: Instance Class Understanding
The task evaluates LLMs’ ability to classify individuals by their correct ontology class.
Each ontology is processed by: 1. Loading the ontology and extracting all classes and instances. 2. For each instance, selecting one asserted class as the correct answer. 3. Generating distractor classes by excluding the target’s ancestors and descendants. 4. Creating multiple-choice questions asking which class the instance belongs to.
Table 9: The instruction and an example of Task U4 Instance Class Understanding.
Instruction: Given a multiple-choice question with options, select the one correct class and output only its letter.
Question: In the ontology of Software, which of the following classes does ’PerkinElmer’ belong to?
A. xps
B. obsolete_MAGE tab format
C. ICS-501 (version 2.3) Image Capture System
D. organization
Answer: D
Task U5: Instance Definition Understanding
The task assesses LLMs’ ability to match instances to their correct textual definitions.
Each ontology is processed by: 1. Loading the ontology and identifying instances with explicit definitions. 2. For each instance, retrieving its definition as the correct answer. 3. Selecting three distractor definitions from other instances of the same classes (or sibling classes). 4. Generating multiple-choice questions asking which definition best describes the given instance.
Table 10: The instruction and an example of Task U5 Instance Definition Understanding.
Instruction: Given a multiple-choice question with options, select the one correct definition and output only its letter.
Question: In the ontology of Semantic Web for Earth and Environmental Terminology, which of the following definitions best describes the instance ’miscible’?
A. Motions describing an equlibrium between pressure gradient force, Coriolis force and the turbulent drag force.
B. Said of two or more liquids that are mutually soluble (i.e. they will dissolve in each other) (McGraw-Hill, 1974). The chemical property of two or more phases that, when brought together, have the ability to mix and form one phase (after AGI, 1980).
C. A system or flow that evolves slowly in time compared to the rotation period of the earth, has a length scale of the deformation radius or larger, and undergoes only limited vertical excursions.
D. The Af climate classification, also known as the tropical rain forest climate classification, is characterized by monthly precipitation in excess of 6 cm all months out of year.
Answer: B
Task R1: Inferred Relation Reasoning
The task evaluates LLMs’ ability to understand inferred ontological relations, which is not explicitly defined in the ontologies.
Each ontology is processed by: 1. Loading the ontology and applying logical reasoning (where feasible) to infer subclass, equivalence, and property-based relations. 2. Extracting explicit and inferred triples using Ontology reasoner (HermiT). 3. Sampling inferred-only triples (i.e., not explicitly stated in the ontology). 4. For each inferred triple, constructing a multiple-choice question with three distractors from non-related classes.
Table 11: The instruction and an example of Task R1 Inferred Relation Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct definition and output only its letter.
Question: In the ontology of Crop, which of the following classes is the inferred superclass of ’Malt friability’?
A. GDONC_M_ppm
B. Barley trait
C. Powdery Mildew susceptibility - Estimation
D. Leaf Stripe incidence
Answer: B
Task R2: Constraint Reasoning
The task assesses LLMs’ understanding of inferred semantic constraints in ontologies, including property ranges, property chains, and existential restrictions.
[…]
Table 12: The instruction and an example of Task R2 Constraint Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct class and output only its letter.
Question: In the ontology of NCI Thesaurus, given the property chain ‘Disease_Has_Finding $\circ$ Disease_Has_Molecular_Abnormality’: Disease_Has_Finding: ‘Disease, Disorder or Finding’ $\to$ ‘Disease, Disorder or Finding’, Disease_Has_Molecular_Abnormality: ‘Disease, Disorder or Finding’ $\to$ ‘Molecular Abnormality’, what class does the chain point to when applied to ‘Finding’ (a subclass of ‘Disease, Disorder or Finding’)?
A. Gene Product
B. Molecular Abnormality
C. Renal Cell Carcinoma Associated with Neuroblastoma
D. Disease, Disorder or Finding
Answer: B
Task R3: Instance Class Reasoning
The task evaluates LLMs’ ability to infer an instance’s class membership via subclass reasoning.
Each ontology is processed by: 1. Extracting explicitly asserted (instance, class) pairs. 2. Inferring additional instance-class pairs via superclass inheritance. 3. Filtering out explicitly stated types to isolate inferred-only instance-class links. 4. Generating multiple-choice questions that ask which class an instance belongs to, with distractors drawn from unrelated classes.
Table 13: The instruction and an example of Task R3 Instance Class Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct inferred class and output only its letter.
Question: In the ontology of Gender, Sex, and Sexual Orientation, which of the following classes does the instance ’Pajubá’ belong to?
A. multi-organism behavior
B. X0 sex-determination system
C. ze/mer/zemself pronouns
D. leather subculture
Answer: A
Task R4: SWRL-Based Logic Reasoning
The task evaluates LLMs’ ability to perform inference over SWRL-style logical rules encoded in OWL ontologies.
Each ontology is processed by: 1. Parsing class-level logical definitions (e.g., existential restrictions, conjunctions) to extract SWRL-like rules. 2. Translating each rule into a natural language question: given a set of conditions (rule body), which logical conclusion (rule head) follows. 3. Generating plausible distractors by altering predicates, argument order, or combining irrelevant rules. 4. Supporting both single-rule and composite multi-rule inference.
Table 14: The instruction and an example of Task R4 SWRL-Based Logic Reasoning.
Instruction: Given a multiple-choice question with options, select the one correct expression and output only its letter.
Question: In the ontology of Financial Industry Business, suppose an individual x satisfies: credit card product(x). Which conclusion is inferred?
A. consumer credit card agreement(x, y)
B. card verification value(x, y)
C. has credit card network(x, y)
D. has credit card network(y, x)
Answer: A
Task R5: Description Logic Reasoning
The task evaluates LLMs’ ability to assess the satisfiability of complex Description Logic (DL) expressions.
Each ontology is processed by: 1. Sampling and synthesizing class expressions using DL constructors (e.g., $\exists$, $\forall$, $\sqcap$, $\geq$, $\neg$). 2. Injecting expressions into a temporary ontology as equivalence axioms and reasoning to check satisfiability. 3. Translating OWL expressions to labeled Description Logic (DL) notation. 4. Filtering expressions by a minimum complexity threshold and balancing between satisfiable and unsatisfiable cases.
Table 15: The instruction and an example of Task R5 Description Logic Reasoning.
Instruction: Given a true-or-false question, determine the correctness of the expression and output only true or false.
Question: Given the following Description Logic expression from the Financial Industry Business Ontology, determine whether it is satisfiable (true or false).
[…]
Answer: False
Task L1: Class Definition Generation
The task evaluates LLMs’ ability to generate or recognize class-level definitions in ontologies.
Each ontology is processed by: 1. Extracting named classes with explicit definitions. 2. Sampling classes based on structural metrics (e.g., depth, siblings, subclass count) to ensure diversity.
Table 16: The instruction and an example of Task L1 Class Definition Generation.
Instruction: Given an open-ended question, please generate a comprehensive text.
Question: Please provide the definition of the concept ’chemist’, as an element in Occupation Ontology.
Answer: Conduct qualitative and quantitative chemical analyses or experiments in laboratories for quality or process control or to develop new products or knowledge.
Task L2: Class Hierarchy Construction
This task assesses LLMs’ ability to reconstruct ontology hierarchies and relations from a set of related classes and their textual definitions.
For each ontology: 1. we select multiple sets of semantically related classes (via subclass/superclass relations). 2. For each set, it extracts a local ontology subgraph (typically 8–15 classes deep) including: subclass axioms, object property and data property triples and class-level annotations (e.g., definitions, comments). 3. Each task presents only the class labels and optional definitions (without explicit structure).
Table 17: The instruction and an example of Task L2 Class Hierarchy Construction.
Instruction: Construct the ontology’s structure and output only the triples—comma-separated—in the form (subject, predicate, object).
Question: Here is a list of classes in the ontology of comic book. Please provide all relationships between classes in triple format.
Classes:
- Agent, - Library, - Museum, - Organization, - OrganizationalUnit, - Partenariat, - Person, - Player, - Publisher, - listened, - music artist, - performer, - sound engineer
Answer:
(Library, subClassOf, Organization)
(Museum, subClassOf, Organization)
(Organization, subClassOf, Agent)
(Organization, subClassOf, Agent)
(OrganizationalUnit, subClassOf, Organization)
(Partenariat, subClassOf, Organization)
(Person, subClassOf, Agent)
(Player, subClassOf, Agent)
(Publisher, subClassOf, Organization)
(listened, subClassOf, Agent)
(music artist, subClassOf, Agent)
(organización formal, subClassOf, Organization)
(performer, subClassOf, Agent)
(sound engineer, subClassOf, Agent)
Task L3: Property and Relation Construction
This task assesses LLMs’ ability to reconstruct ontology property structures from a set of interrelated classes and a list of available property names.
For each ontology: 1. we select multiple sets of structurally connected classes (based on subclass and property usage patterns). 2. For each set, we extract a local ontology subgraph (8–15 classes) that includes object property and data property assertions among the classes, as well as relevant annotations. 3. Each task presents only the class labels (with optional comments) and a list of used object/data property names—without revealing any relational structure—requiring models to recover all valid property triples and their characteristics (e.g., functional, symmetric, transitive).
Table 18: The instruction and an example of Task L3 Property and Relation Construction.
Instruction: Construct the ontology’s structure and output only the triples—comma-separated—in the form (subject, predicate, object).
Question: Here is a list of classes and properties in the ontology of Financial Industry Business. Please provide all relationships properties in triple format. Classes: - announcement, - context, - date, - debt underwriting closing, - dissemination, - issuance closing, - issuance process activity, - issuance settlement, - jurisdiction, - m t n registration, - offering, - registration, - underwriter takedown, - underwriting issuance closing, - underwriting process activity
Object Properties: - actual closing date, - facililtates remarketing, - formal award date and time, - has agent, - has bond counsel, - has financial advisor, - has issuer counsel, - has remarketing agent, - has trustee, - is bond counsel, - issuance settlement date, - makes decision on, - maximum over allotment percentage, - obligor to, - prints, - refers to, - registered under, - step, - subscribes to, - underwritten by
Data Properties: - announcement comments, - maximum over allotment shares, - over allotment available, - over allotment expiry period days, - sale method
Answer:
(issuance closing, actual closing date, date)
(issuance settlement, issuance settlement date, date)
(registration, registered under, jurisdiction)
(underwriting issuance closing, refers to, underwriter takedown)
Task L4: Constraint Construction
This task assesses LLMs’ ability to recover property-level constraints from an ontology.
For each ontology: 1. We sample multiple sets of ontology properties (object and data properties), ensuring each set is sufficiently rich. 2. For each set, we extract a local subgraph of related classes by traversing the domain and range of the selected properties, expanding outward (via subclass relations) to reach a representative size. 3. Each task presents a list of class labels (with optional definitions/comments) and a list of property names.
Table 19: The instruction and an example of Task L4 Constraint Construction.
Instruction: Construct the ontology’s structure (constraint) and generate only the triples—comma-separated—in the form (subject, domain/range, object).
Question: Here is a list of classes and properties in the ontology of Semantic Web for Earth and Environmental Terminology. Please generate the property constraints (domain or range) in triple form. Classes: - chemical quantity, - coordinates, - electromagnetic quantity, - electromagnetic radiation quantity, - extensive property, - fluence, - intensive property, - minimum, - number density, - physical quantity, - quantity, - scientific quantity, - species, - thermodynamic quantity, - unit
Properties: - nonoverlapping, - kills, - has minimum, - has coordinates, - has upper boundary, - has average duration, - has unit, - impacts, - causes, - due to
Answer:
(has coordinates, range, coordinates)
(has minimum, range, minimum)
(has unit, domain, quantity)
(has unit, range, unit)
(kills, range, species)
Task L5: Ontology Alignment
This task assesses LLMs’ ability to align classes, properties, and instances between two ontologies. It is derived from the VeeAlign [^4] https://github.com/Remorax/VeeAlign framework [42], a deep learning-based ontology alignment model that leverages a dual-attention mechanism to capture both syntactic and semantic information of entities. We construct this task by directly extracting input-output pairs from the VeeAlign dataset, where ontology entity pairs serve as inputs and their corresponding alignment labels are treated as answers. This setup enables evaluation of LLMs’ understanding of ontological structures and their capacity for semantic alignment across domains and languages.
Table 20: The instruction and an example of Task L5 Ontology Alignment.
Instruction: Your task is to perform ontology alignment. Please only generate the matched pairs as tuples in the format (entity1, entity2).
Question: Please align the classes between these two ontologies:
- Ontology 1 classes: ’Person’, ’Review expertise’, ’Reviewer’, ’Regular author’, ’Co-chair’, ’Committee’, ’Committee member’, ’Abstract’, ’Extended abstract’, ’Important dates’, ’Invited speaker’, ’Conference document’, ’Organization’, ’Regular contribution’, ’Poster’, ’Written contribution’, ’Conference applicant’, ’Program committee’, ’Organizing committee’, ’Steering committee’, ’Conference volume’, ’Conference’, ’Call for paper’, ’Paper’, ’Invited talk’, ’Conference part’, ’Presentation’, ’Chair’, ’Track’, ’Workshop’, ’Tutorial’, ’Review’, ’Topic’, ’Late paid applicant’, ’Paid applicant’, ’Early paid applicant’, ’Organizer’, ’Publisher’, ’Track-workshop chair’, ’Conference www’, ’Conference fees’, ’Review preference’
- Ontology 2 classes: ’SlideSet’, ’Review’, ’Programme’, ’Paper’, ’MealMenu’, ’Document’, ’Topic’, ’ConferenceEvent’, ’Call’, ’ReviewRating’, ’Organization’, ’ReviewForm’, ’Conference’, ’PersonalHistory’, ’Sponsorship’, ’ReviewQuestion’, ’Person’, ’Country’, ’ConferenceSession’, ’ContactInformation’, ’Place’, ’WeekRejectRating’, ’AcceptRating’, ’AccpetIfRoomRating’, ’RejectRating’, ’MealEvent’, ’NonAcademicEvent’, ’BreakEvent’, ’CADTopic’, ’MultimediaTopic’, ’CommunicationsTopic’, ’AntennasTopic’, ’PerformanceTopic’, ’TestOnlyTopic’, ’MedicineTopic’, ’SecurityTopic’, ’MobileComputingTopic’, ’CryptographyTopic’, ’ConferenceChair’, ’Author’, ’DiningPlace’, ’Presenter’, ’AcademiaOrganization’, ’IndustryOrganization’, ’NGO’, ’ActivePaper’, ’PendingPaper’, ’WithdrawnPaper’, ’RejectedPaper’, ’AcceptedPaper’, ’PublishedPaper’, ’Attendee’, ’CoffeeBreak’, ’FreeTimeBreak’, ’MealBreak’, ’Reviewer’, ’ClosingTalk’, ’WelcomeTalk’, ’TalkEvent’, ’SessionChair’, ’TPCMember’, ’TwoLevelConference’, ’CallForPapers’, ’PaperPresentation’, ’Workshop’, ’AcademicEvent’, ’Excursion’, ’SocialEvent’, ’CallForReviews’, ’ConferenceVenuePlace’, ’CallForManuscripts’, ’Reception’, ’MeetingRoomPlace’, ’AccommodationPlace’, ’RatedPapers’, ’TravelGrant’, ’ConferenceDinner’
Answer: (’Person’, ’Person’)
(’Organization’, ’Organization’)
(’Reviewer’, ’Reviewer’)
(’has the first name’, ’hasFirstName’)
(’Conference part’, ’ConferenceEvent’)
(’Workshop’, ’Workshop’)
(’Conference document’, ’Document’)
(’Paper’, ’Paper’)
(’has the last name’, ’hasLastName’)
(’Review’, ’Review’)
(’Conference volume’, ’Conference’)
(’Topic’, ’Topic’)
(’Regular author’, ’Author’)
Appendix E Expert Verification
Guidelines
@@ Appendix E Expert Verification
Appendix F Details of Experiments
@@
The models and their access are shown in table 24.
Table 24: Models and its repositories.
@@
Medicine-hat-7B https://huggingface.co/AdaptLLM/medicine-chat
@@
Appendix G Additional Evaluation Result
In this section, we present the full results, including zero-shot, two-shot and four-shot settings in Table 25.
Table 25: All performance of LLMs under zero-, two- and four-shot settings.
@@ 3-4B Models
Qwen2.5-3B Zero 77.8 86.3 80.3 85.8 77.8 81.5 74.9 65.7 62.5 67.9 5.9 0.1 0.0 0.2 6.7
Two 81.6 88.3 76.3 86.9 79.2 79.4 91.4 63.7 68.0 68.7 7.5 8.5 0.0 1.0 6.5
Four 81.3 87.7 81.3 89.5 80.7 79.7 91.4 63.6 70.0 68.7 8.1 8.2 0.0 1.8 4.4
Phi4-4B Zero 77.5 91.1 75.5 87.2 78.8 80.2 80.7 63.5 59.1 51.1 9.4 0.1 0.0 0.0 0.1
Two 79.5 92.4 76.3 87.4 77.0 83.0 91.8 61.3 68.4 64.3 5.3 0.3 0.1 0.2 2.1
Four 78.9 93.2 81.3 87.6 78.9 82.5 91.3 64.7 70.9 63.6 6.1 0.3 0.2 0.9 1.7
@@ 7-8B Models
Qwen2.5-7B Zero 83.1 90.6 77.6 90.1 83.6 87.6 88.2 73.9 66.0 68.6 5.6 0.4 0.1 0.3 16.2
Two 83.7 93.3 77.1 89.9 82.5 83.7 94.9 67.5 71.7 69.5 10.4 8.8 0.3 1.7 15.3
Four 83.8 93.1 76.3 90.8 83.6 82.0 95.3 67.7 73.8 69.0 16.9 15.0 0.6 2.0 10.2
Ollm-wiki-7B Zero 74.3 84.5 67.2 81.4 77.0 65.2 83.3 53.4 57.3 9.3 6.0 0.1 0.0 0.2 0.1
Two 77.5 87.6 71.2 85.9 75.0 64.9 86.9 53.1 68.0 64.0 5.1 13.3 0.2 1.2 3.3
Four 76.5 86.1 74.1 85.3 77.7 67.0 89.9 55.8 68.4 56.1 4.9 16.5 0.5 1.5 3.4
Ollm-arxiv-7B Zero 74.1 84.4 67.5 81.5 77.0 64.2 82.8 53.1 56.6 8.4 6.0 0.1 0.0 0.0 8.3
Two 77.4 87.5 70.4 85.9 74.9 64.0 87.0 52.9 67.9 62.4 5.1 13.5 0.2 1.6 3.2
Four 76.7 85.8 73.6 84.9 77.4 66.3 89.8 55.7 68.4 55.9 4.9 16.1 0.5 1.8 3.8
Ministral-8B Zero 78.9 88.8 62.4 83.9 79.5 81.0 88.5 60.1 62.4 52.7 10.3 0.1 0.0 0.1 16.4
Two 81.5 93.6 68.3 89.8 80.4 81.7 89.0 61.2 67.7 68.6 25.1 23.2 0.1 1.9 10.2
Four 81.9 93.7 96.5 89.1 81.8 80.0 90.0 63.3 71.4 70.1 26.6 22.6 0.3 2.2 9.8
Llama3.1-8B Zero 79.8 87.2 74.9 88.4 81.1 79.8 84.2 72.3 62.2 68.9 6.4 0.1 0.0 0.1 15.3
Two 82.6 91.5 70.4 89.9 81.9 85.2 92.5 64.6 69.4 68.3 4.5 13.3 0.1 2.5 10.5
Four 83.2 91.5 78.7 90.8 83.1 84.5 93.4 5.9 75.1 68.0 4.4 15.9 0.4 2.1 9.7
Internlm3-8B Zero 83.1 90.9 72.0 88.9 82.4 88.5 90.5 73.8 67.2 62.9 6.3 0.2 0.0 0.4 12.0
Two 83.8 91.7 71.2 88.8 82.3 75.3 94.4 62.0 72.9 67.4 5.4 13.0 0.0 1.1 11.8
Four 84.1 92.9 82.1 87.8 82.5 72.5 95.1 61.6 77.5 68.6 6.4 15.5 0.2 1.4 8.4
Aya-8B Zero 77.1 85.8 62.4 83.8 77.9 73.0 78.0 62.6 57.4 63.6 7.5 0.1 0.0 0.0 6.3
Two 77.4 88.6 69.6 88.0 77.8 73.4 79.5 59.7 64.4 67.9 4.5 11.5 0.0 1.1 7.2
Four 76.8 88.4 72.5 89.3 78.5 75.3 81.6 63.4 68.8 68.8 4.4 11.8 0.3 1.2 5.0
@@
Qwen2.5-14B Zero 86.6 92.0 75.5 91.4 85.8 89.6 94.0 76.4 71.2 63.6 5.6 0.1 0.1 1.0 19.5
Two 85.6 93.4 78.9 91.3 85.8 89.1 95.9 71.2 78.6 69.0 5.3 18.3 0.3 2.0 18.1
Four 86.9 94.5 82.4 92.5 87.3 87.2 96.4 74.6 81.9 68.7 5.3 19.1 0.5 3.8 16.8
Mistral-22B Zero 83.9 90.4 69.6 88.6 84.4 86.3 86.9 69.3 64.0 54.3 7.5 0.1 0.0 0.8 15.8
Two 87.0 94.5 74.4 89.3 86.4 88.6 94.2 67.0 78.8 68.9 22.2 18.3 0.2 3.0 13.5
Four 86.9 95.2 79.7 89.8 87.2 89.0 95.6 69.1 80.6 69.3 21.9 0.7 0.2 2.5 5.1
Qwen2.5-32B Zero 88.0 90.6 81.9 91.2 87.2 89.7 95.5 76.8 72.4 68.4 5.7 1.6 0.1 1.5 20.3
Two 88.9 94.7 84.8 91.2 88.7 88.2 96.6 76.0 81.8 69.2 5.3 18.9 0.3 2.9 22.6
Four 88.9 94.7 85.3 91.8 89.3 87.9 97.0 79.0 84.1 68.9 5.3 19.4 0.7 2.8 19.1
QwQ-32B Zero 82.2 89.6 77.1 88.9 81.5 84.0 92.5 70.8 60.6 63.4 6.1 1.1 0.2 1.0 18.0
Two 88.1 94.4 82.9 89.7 87.9 87.8 95.9 71.8 82.0 58.7 7.7 14.9 0.1 1.9 19.4
Four 88.0 95.1 86.1 91.0 88.8 86.9 96.4 75.4 84.0 68.9 8.9 4.9 0.4 4.0 16.7
Aya-32B Zero 81.2 90.5 61.6 89.7 82.3 85.5 83.1 70.3 66.0 68.8 6.5 0.1 0.1 0.5 19.3
Two 85.4 94.9 74.4 91.1 85.0 80.4 92.3 65.5 75.0 70.7 4.6 17.5 0.1 4.6 14.4
Four 85.5 95.1 68.5 90.2 85.8 78.2 94.3 68.8 78.6 71.6 4.6 18.8 0.4 6.1 14.5

### v3

@@ Appendix B Task Statistics
Appendix C Expert Verification
@@
Appendix D Details of Experiments
@@
The models and their repositories are available in Table 9.
@@
Medicine-chat-7B https://huggingface.co/AdaptLLM/medicine-chat
@@
Table 9: Models and its repositories.
Appendix E Additional Evaluation Result
All the results for few-shot experiments and chain-of-thought experiments are shown in Table 10 and Table 11, respectively.
@@ 3-4B Models
Qwen2.5-3B Zero 77.8 86.3 80.3 85.8 77.8 81.5 74.9 65.7 62.5 67.9 79.8 0.1 0.0 0.2 6.7
Two 81.6 88.3 76.3 86.9 79.2 79.4 91.4 63.7 68.0 68.7 80.5 8.5 0.0 1.0 6.5
Four 81.3 87.7 81.3 89.5 80.7 79.7 91.4 63.6 70.0 68.7 81.0 8.2 0.0 1.8 4.4
Phi4-4B Zero 77.5 91.1 75.5 87.2 78.8 80.2 80.7 63.5 59.1 51.1 82.4 0.1 0.0 0.0 0.1
Two 79.5 92.4 76.3 87.4 77.0 83.0 91.8 61.3 68.4 64.3 83.1 0.3 0.1 0.2 2.1
Four 78.9 93.2 81.3 87.6 78.9 82.5 91.3 64.7 70.9 63.6 83.6 0.3 0.2 0.9 1.7
@@ 7-8B Models
Qwen2.5-7B Zero 83.1 90.6 77.6 90.1 83.6 87.6 88.2 73.9 66.0 68.6 79.8 0.4 0.1 0.3 16.2
Two 83.7 93.3 77.1 89.9 82.5 83.7 94.9 67.5 71.7 69.5 80.6 8.8 0.3 1.7 15.3
Four 83.8 93.1 76.3 90.8 83.6 82.0 95.3 67.7 73.8 69.0 81.1 15.0 0.6 2.0 10.2
Ollm-wiki-7B Zero 74.3 84.5 67.2 81.4 77.0 65.2 83.3 53.4 57.3 9.3 79.0 0.1 0.0 0.2 0.1
Two 77.5 87.6 71.2 85.9 75.0 64.9 86.9 53.1 68.0 64.0 79.8 13.3 0.2 1.2 3.3
Four 76.5 86.1 74.1 85.3 77.7 67.0 89.9 55.8 68.4 56.1 80.3 16.5 0.5 1.5 3.4
Ollm-arxiv-7B Zero 74.1 84.4 67.5 81.5 77.0 64.2 82.8 53.1 56.6 8.4 79.1 0.1 0.0 0.0 8.3
Two 77.4 87.5 70.4 85.9 74.9 64.0 87.0 52.9 67.9 62.4 79.9 13.5 0.2 1.6 3.2
Four 76.7 85.8 73.6 84.9 77.4 66.3 89.8 55.7 68.4 55.9 80.4 16.1 0.5 1.8 3.8
Ministral-8B Zero 78.9 88.8 62.4 83.9 79.5 81.0 88.5 60.1 62.4 52.7 82.6 0.1 0.0 0.1 16.4
Two 81.5 93.6 68.3 89.8 80.4 81.7 89.0 61.2 67.7 68.6 83.4 23.2 0.1 1.9 10.2
Four 81.9 93.7 96.5 89.1 81.8 80.0 90.0 63.3 71.4 70.1 83.9 22.6 0.3 2.2 9.8
Llama3.1-8B Zero 79.8 87.2 74.9 88.4 81.1 79.8 84.2 72.3 62.2 68.9 79.4 0.1 0.0 0.1 15.3
Two 82.6 91.5 70.4 89.9 81.9 85.2 92.5 64.6 69.4 68.3 80.2 13.3 0.1 2.5 10.5
Four 83.2 91.5 78.7 90.8 83.1 84.5 93.4 5.9 75.1 68.0 80.7 15.9 0.4 2.1 9.7
Internlm3-8B Zero 83.1 90.9 72.0 88.9 82.4 88.5 90.5 73.8 67.2 62.9 79.7 0.2 0.0 0.4 12.0
Two 83.8 91.7 71.2 88.8 82.3 75.3 94.4 62.0 72.9 67.4 80.5 13.0 0.0 1.1 11.8
Four 84.1 92.9 82.1 87.8 82.5 72.5 95.1 61.6 77.5 68.6 81.0 15.5 0.2 1.4 8.4
Aya-8B Zero 77.1 85.8 62.4 83.8 77.9 73.0 78.0 62.6 57.4 63.6 80.5 0.1 0.0 0.0 6.3
Two 77.4 88.6 69.6 88.0 77.8 73.4 79.5 59.7 64.4 67.9 81.3 11.5 0.0 1.1 7.2
Four 76.8 88.4 72.5 89.3 78.5 75.3 81.6 63.4 68.8 68.8 81.8 11.8 0.3 1.2 5.0
@@
Qwen2.5-14B Zero 86.6 92.0 75.5 91.4 85.8 89.6 94.0 76.4 71.2 63.6 79.9 0.1 0.1 1.0 19.5
Two 85.6 93.4 78.9 91.3 85.8 89.1 95.9 71.2 78.6 69.0 80.8 18.3 0.3 2.0 18.1
Four 86.9 94.5 82.4 92.5 87.3 87.2 96.4 74.6 81.9 68.7 81.3 19.1 0.5 3.8 16.8
Mistral-22B Zero 83.9 90.4 69.6 88.6 84.4 86.3 86.9 69.3 64.0 54.3 80.1 0.1 0.0 0.8 15.8
Two 87.0 94.5 74.4 89.3 86.4 88.6 94.2 67.0 78.8 68.9 81.0 18.3 0.2 3.0 13.5
Four 86.9 95.2 79.7 89.8 87.2 89.0 95.6 69.1 80.6 69.3 81.5 0.7 0.2 2.5 5.1
Qwen2.5-32B Zero 88.0 90.6 81.9 91.2 87.2 89.7 95.5 76.8 72.4 68.4 80.0 1.6 0.1 1.5 20.3
Two 88.9 94.7 84.8 91.2 88.7 88.2 96.6 76.0 81.8 69.2 80.9 18.9 0.3 2.9 22.6
Four 88.9 94.7 85.3 91.8 89.3 87.9 97.0 79.0 84.1 68.9 81.4 19.4 0.7 2.8 19.1
QwQ-32B Zero 82.2 89.6 77.1 88.9 81.5 84.0 92.5 70.8 60.6 63.4 79.4 1.1 0.2 1.0 18.0
Two 88.1 94.4 82.9 89.7 87.9 87.8 95.9 71.8 82.0 58.7 80.3 14.9 0.1 1.9 19.4
Four 88.0 95.1 86.1 91.0 88.8 86.9 96.4 75.4 84.0 68.9 80.8 4.9 0.4 4.0 16.7
Aya-32B Zero 81.2 90.5 61.6 89.7 82.3 85.5 83.1 70.3 66.0 68.8 79.3 0.1 0.1 0.5 19.3
Two 85.4 94.9 74.4 91.1 85.0 80.4 92.3 65.5 75.0 70.7 80.2 17.5 0.1 4.6 14.4
Four 85.5 95.1 68.5 90.2 85.8 78.2 94.3 68.8 78.6 71.6 80.7 18.8 0.4 6.1 14.5